# TESLA-X: An effective method to search for sub-threshold lensed gravitational waves with a targeted population model

### v2

@@ Appendix A Choosing a suitable pair of parameters to construct the Gaussian KDE
In the main text, we construct the Gaussian KDE in the chirp mass $\mathcal{M}_{c}$ effective spin $\chi_{\text{eff}}$ parameter space. It has also been suggested that the component masses $m_{1}$-$m_{2}$ parameter space may be a better choice. In this appendix, we present some preliminary investigation results for this question.
@@
In this subsection, we redo the TESLA-X analysis with the Gaussian KDE constructed in the component masses $m_{1}$-$m_{2}$ parameter space instead of the chirp mass $\mathcal{M}_{c}$ effective spin $\chi_{\text{eff}}$ parameter space.
@@ A.1.1 Constructing the TESLA-X bank and population model
Figure 11: The contour map of the Gaussian Kernel Density Estimation function obtained for MS220508a in the $m_{1}$ $m_{2}$ space. The colors represent the probability density estimated by the KDE function.
We again take the lowest contour level ($0.00002$) as a boundary, and use all the templates that fall within this boundary to construct a TESLA-X targeted template bank in the $m_{1}$-$m_{2}$ space. For simplicity, we will call the TESLA-X bank created based on the KDE constructed in the $m_{1}$-$m_{2}$ space as the “TESLA-X (component mass) bank”, and the TESLA-X bank in the main text that is created based on the KDE constructed in the $\mathcal{M}_{c}$ $\chi_{\text{eff}}$ space as the “TESLA-X (main) bank”. Figure 12 shows the templates in the full template bank (in grey) and in the “TESLA-X (component mass) bank” (in red) in the $m_{1}$-$m_{2}$ space. The yellow curve marks the lowest contour level we obtained for the estimated Gaussian KDE function we showed in Figure 11. There are $159043$ templates in the “TESLA-X (component mass) bank”.
@@
Similar to the main text, we use the estimated KDE function for MS220508a and the “TESLA-X (component mass) bank” to construct a targeted population model, following the same steps detailed in IV and Fong (2018).
@@
To compare the effectiveness of the “TESLA-X (main) bank” and “TESLA-X (component mass) bank” for the case of MS220508a, we perform an additional search using the “TESLA-X (component mass) bank” together with the targeted population model over the same stretch of mock data described in the main text to try to recover the same set of lensed injections used in the injection campaign. Table 3 summarizes the findings. We can see that “TESLA-X (component mass)” is performing even worse than the full template bank, with a reduction in recovered lensed injections by $-0.3\%$. This should not be surprising, however, given that the “ TESLA-X (component mass) bank” has more templates than the “TESLA-X (main) bank” (almost six times more templates). This results in a much higher trials factor and, hence, a larger noise background, making it harder to recover the simulated lensed injections.
@@
Figure 13: (Top panel) The sensitive range v.s. FAR for MS220508a-alike signals using the full template bank (black), “TESLA-X (main) bank” (red), and “TESLA-X (component mass) bank” (blue) respectively. The shaded band for each curve represents the corresponding $1$-sigma region. (Bottom panel) The corresponding percentage changes in sensitive range v.s. FAR for the different banks. We note that while the “TESLA-X (main) bank” improves in terms of sensitivity towards MS220508a-alike (lensed) signals, “TESLA-X (component mass) bank” is performing worse than the full template bank, mainly due to the significant increase in trials factors caused by the large number of templates.
Furthermore, if we look at the sensitive range v.s. FAR for MS220508a-alike signals using the full template bank, “TESLA-X (main) bank” and “TESLA-X (component mass) bank” respectively (see Figure 13), we can see that the “TESLA-X (component mass) bank” is performing even worse than the full template bank. Therefore, in the case of MS220508a, chirp mass $\mathcal{M}_{c}$ effective spin $\chi_{\text{eff}}$ would be the better pair of parameters to construct the Gaussian KDE in the TESLA-X method.
@@ A.2 Analyzing another mock event MS220510ae with the TESLA-X method
Figure 14: The contour map of the Gaussian Kernel Density Estimation function obtained for MS220510ae in the $\mathcal{M}_{c}$ $\chi_{\text{eff}}$ space. The colors represent the probability density estimated by the KDE function.
Figure 15: The contour map of the Gaussian Kernel Density Estimation function obtained for MS220510ae in the $m_{1}$ $m_{2}$ space. The colors represent the probability density estimated by the KDE function.
@@
Figure 16: (Top panel) The sensitive range v.s. FAR for MS220510ae-alike signals using the full template bank (black), “TESLA-X (main) bank” (red), and “TESLA-X (component mass) bank” (blue) respectively. The shaded band for each curve represents the corresponding $1$-sigma region. (Bottom panel) The corresponding percentage changes in sensitive range v.s. FAR for the different banks. We note that, unlike the case of MS220508a in the main text, the “TESLA-X (component mass) bank” improves in terms of sensitivity towards MS220510ae-alike (lensed) signals, “TESLA-X (main) bank” performs worse than the full template bank. This is mainly because spins are typically poorly measured for high chirp mass systems.
@@ A.3 Conclusion
From the two investigations, we note that “chirp mass - effective spins” is a better parameter pair to be used in constructing the Gaussian KDE for the TESLA-X method, if the target event is a low chirp mass system. On the other hand, component masses “mass $1$ mass $2$” will be a better parameter pair to be used in constructing the Gaussian KDE for the TESLA-X method, if the target event is a high chirp mass system. We note that these findings are entirely expected: While chirp masses and spins are well-measured in low-mass systems, they are not well-measured in high-mass systems since their signals have much fewer in-band cycles. As future work, we will investigate how we can improve the construction of the Gaussian KDE functions for the TESLA-X method. For instance, we can see that the Gaussian KDEs can exceed the physical boundary of the original template bank (See the yellow boundary in 12). Regions near the physical boundary of the original template bank may require special care. We may also explore the idea of first constructing and applying a separate KDE to “flatten” the TESLA-X bank (i.e., to construct a uniform mass model) before applying the Gaussian KDE representing the targeted population model.

### v5

@@ Appendix A Choosing a suitable pair of parameters to construct the Gaussian KDE
In the main text, we construct the Gaussian KDE in the chirp mass $\mathcal{M}_{c}$ - effective spin $\chi_{\text{eff}}$ parameter space. It has also been suggested that the component masses $m_{1}$-$m_{2}$ parameter space may be a better choice. In this appendix, we present some preliminary investigation results for this question.
@@
In this subsection, we redo the TESLA-X analysis with the Gaussian KDE constructed in the component masses $m_{1}$-$m_{2}$ parameter space instead of the chirp mass $\mathcal{M}_{c}$ - effective spin $\chi_{\text{eff}}$ parameter space.
@@ A.1.1 Constructing the TESLA-X bank and population model
Figure 11: The contour map of the Gaussian Kernel Density Estimation function obtained for MS220508a in the $m_{1}$ - $m_{2}$ space. The colors represent the probability density estimated by the KDE function.
We again take the lowest contour level ($0.00002$) as a boundary, and use all the templates that fall within this boundary to construct a TESLA-X targeted template bank in the $m_{1}$-$m_{2}$ space. For simplicity, we will call the TESLA-X bank created based on the KDE constructed in the $m_{1}$-$m_{2}$ space as the “TESLA-X (component mass) bank”, and the TESLA-X bank in the main text that is created based on the KDE constructed in the $\mathcal{M}_{c}$ - $\chi_{\text{eff}}$ space as the “TESLA-X (main) bank”. Figure 12 shows the templates in the full template bank (in grey) and in the “TESLA-X (component mass) bank” (in red) in the $m_{1}$-$m_{2}$ space. The yellow curve marks the lowest contour level we obtained for the estimated Gaussian KDE function we showed in Figure 11. There are $159043$ templates in the “TESLA-X (component mass) bank”.
@@
Similar to the main text, we use the estimated KDE function for MS220508a and the “TESLA-X (component mass) bank” to construct a targeted population model, following the same steps detailed in 4 and (Fong, 2018).
@@
To compare the effectiveness of the “TESLA-X (main) bank” and “TESLA-X (component mass) bank” for the case of MS220508a, we perform an additional search using the “TESLA-X (component mass) bank” together with the targeted population model over the same stretch of mock data described in the main text to try to recover the same set of lensed injections used in the injection campaign. Table 3 summarizes the findings. We can see that “TESLA-X (component mass)” is performing even worse than the full template bank, with a reduction in recovered lensed injections by $0.3\%$. This should not be surprising, however, given that the “ TESLA-X (component mass) bank” has more templates than the “TESLA-X (main) bank” (almost six times more templates). This results in a much higher trials factor and, hence, a larger noise background, making it harder to recover the simulated lensed injections.
@@
Figure 13: (Top panel) The sensitive range v.s. FAR for MS220508a-alike signals using the full template bank (black), “TESLA-X (main) bank” (red), and “TESLA-X (component mass) bank” (blue) respectively. The shaded band for each curve represents the corresponding 11-sigma region. (Bottom panel) The corresponding percentage changes in sensitive range v.s. FAR for the different banks. We note that while the “TESLA-X (main) bank” improves in terms of sensitivity towards MS220508a-alike (lensed) signals, “TESLA-X (component mass) bank” is performing worse than the full template bank, mainly due to the significant increase in trials factors caused by the large number of templates.
Furthermore, if we look at the sensitive range v.s. FAR for MS220508a-alike signals using the full template bank, “TESLA-X (main) bank” and “TESLA-X (component mass) bank” respectively (see Figure 13), we can see that the “TESLA-X (component mass) bank” is performing even worse than the full template bank. Therefore, in the case of MS220508a, chirp mass $\mathcal{M}_{c}$ - effective spin $\chi_{\text{eff}}$ would be the better pair of parameters to construct the Gaussian KDE in the TESLA-X method.
@@ A.2 Analyzing another mock event MS220510ae with the TESLA-X method
Figure 14: The contour map of the Gaussian Kernel Density Estimation function obtained for MS220510ae in the $\mathcal{M}_{c}$ - $\chi_{\text{eff}}$ space. The colors represent the probability density estimated by the KDE function.
Figure 15: The contour map of the Gaussian Kernel Density Estimation function obtained for MS220510ae in the $m_{1}$ - $m_{2}$ space. The colors represent the probability density estimated by the KDE function.
@@
Figure 16: (Top panel) The sensitive range v.s. FAR for MS220510ae-alike signals using the full template bank (black), “TESLA-X (main) bank” (red), and “TESLA-X (component mass) bank” (blue) respectively. The shaded band for each curve represents the corresponding 11-sigma region. (Bottom panel) The corresponding percentage changes in sensitive range v.s. FAR for the different banks. We note that, unlike the case of MS220508a in the main text, the “TESLA-X (component mass) bank” improves in terms of sensitivity towards MS220510ae-alike (lensed) signals, “TESLA-X (main) bank” performs worse than the full template bank. This is mainly because spins are typically poorly measured for high chirp mass systems.
@@ A.3 Conclusion
From the two investigations, we note that “chirp mass - effective spins” is a better parameter pair to be used in constructing the Gaussian KDE for the TESLA-X method, if the target event is a low chirp mass system. On the other hand, component masses “mass $1$ - mass 22” will be a better parameter pair to be used in constructing the Gaussian KDE for the TESLA-X method, if the target event is a high chirp mass system. We note that these findings are entirely expected: While chirp masses and spins are well-measured in low-mass systems, they are not well-measured in high-mass systems since their signals have much fewer in-band cycles. As future work, we will investigate how we can improve the construction of the Gaussian KDE functions for the TESLA-X method. For instance, we can see that the Gaussian KDEs can exceed the physical boundary of the original template bank (See the yellow boundary in 12). Regions near the physical boundary of the original template bank may require special care. We may also explore the idea of first constructing and applying a separate KDE to “flatten” the TESLA-X bank (i.e., to construct a uniform mass model) before applying the Gaussian KDE representing the targeted population model.